# One Homography is All You Need: IMM-based Joint Homography and Multiple Object State Estimation

### v3

@@ 5 Experiments
This section describes the experimental procedure used to find optimal hyperparameters for IMM-JHSE and reports the results of experiments on validation data. The datasets considered during these experiments are the MOT17 [35] and DanceTrack [33] datasets. Results on the test sets are reported in Section 6 and include the MOT20 [36] and KITTI [34] datasets.
@@ 5.1 Parameter selection
The largest assumption of IMM-JHSE is that an initial homography estimate is available. For all experiments and test set results, the camera calibration matrices provided in [32] are used – except for the KITTI dataset, which provides its own. Furthermore, the estimates of $\mathbf{\hat{A}}_{t}$ (2) provided by [32] are used for the MOT17 and MOT20 datasets, but for the DanceTrack and KITTI datasets $\mathbf{\hat{A}}_{t}$ is estimated as mentioned in Section 3.
@@ 5.1.1 Filter initialisation
In addition to the parameters above, the covariance matrices with which each filter is initialised must be specified. The initial covariance matrix associated with the ground plane coordinates $\mathbf{x}^{W}_{t}$ is obtained similar to [32]:
@@
This paper uses the pattern search algorithm [83] to find optimum values for the hyperparameters. For each experiment described subsequently, the initial values are set as follows: $\sigma_{x}=5,\sigma_{y}=5,\alpha_{1}=\alpha_{2}=\alpha_{3}=0.5,\Omega=30,b=0,d_{\text{high}}=0.6,d_{\text{low}}=0.5,p_{s,s}=p_{d,d}=p_{W,W}=p_{I,I}=0.9,v=0.5$, and the number of pattern search iterations is restricted to 200. The objective function selected for optimisation is HOTA for the MOT17, MOT20 and KITTI datasets and association accuracy (AssA) for the DanceTrack dataset. Other evaluation metrics include IDF1 and multiple object tracking accuracy (MOTA). The reader is referred to [84] for details regarding these metrics. In all cases, evaluation is performed with the TrackEval library [85]. For the DanceTrack, KITTI and MOT20 datasets, the average of the objective functions of each video in the entire dataset is optimised, while for the MOT17 dataset, the objective function is optimised per video.
The parameters obtained for IMM-JHSE with the pattern search algorithm are given in Table 9.
@@ 5.2 Ablation study on the DanceTrack dataset
Table 1 reports the results of an ablation study on the DanceTrack validation dataset. The second column refers to using the dynamic measurement noise estimation method described in Section 4.3. When dynamic measurement noise is not used (as indicated with X), (16) is used to provide an estimate of the measurement noise at each time step. When the dynamic estimation method is not used, HOTA decreases from $63.81$ to $60.80$. Thus, the dynamic noise estimation method contributes significantly to the success of IMM-JHSE.
@@
UCMC [32] X X X X 60.42 62.64
UCMC+ [32] X X X X 59.18 62.52
Table 1: Ablation study on the DanceTrack validation set, with detections obtained from [11]. The columns denote various components of IMM-JHSE as described in Section 5. UCMC+ refers to the method of [32] with camera motion compensation.
@@ 5.3 Investigating association score mixture methods
Table 2: Validation results on the MOT17 and DanceTrack training sets with detections obtained from [9], where the way in which bounding box information is used is varied.
@@ 5.4 Investigation of the chosen objective function
Table 3: Validation results on the MOT17 and DanceTrack training sets with detections obtained from [9], where the optimised metric is varied. Note that, for some datasets, the results depend greatly on the metric being optimised.
@@ 6 Results
This section presents the results of IMM-JHSE evaluated on the DanceTrack, MOT17, MOT20 and KITTI test datasets.
@@ 6.1 Results on DanceTrack
ByteTrack [9] 47.3 52.5 89.5 31.4
C-BIoU [12] 60.6 61.6 91.6 45.4
OC-SORT [11] 55.1 54.9 92.2 40.4
UCMCTrack [32] 63.4 65.0 88.8 51.1
UCMCTrack+ [32] 63.6 65.0 88.9 51.3
@@
FusionTrack [44] 65.3 73.3 90.1 57.5
MeMOTR [45] 66.7 70.6 94 53
MeMOTR [45] 68.5 71.2 89.9 58.4
MOTIP [46] 67.5 72.2 90.3 57.6
MOTIP [46] 70 75.1 91 60.8
Deformable DETR [86]
DAB-Deformable DETR [87]
Table 4: Results on the DanceTrack test set. For the tracking-by-detection methods, the detections are obtained from [9].
@@ 6.2 Results on the KITTI dataset
At the time of writing, IMM-JHSE is placed twenty-fourth on the KITTI-car leaderboard [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking_detail.php?result=c3e8939f18cba7b23a2e05a92d4df9e994eaaf91. In addition, all methods that are placed above it are 3D MOT methods that make use of regular 3D measurements, except for RAM [88], which uses a self-supervised encoder that optimises a Markov walk along a space-time graph. Table 6 compares IMM-JHSE and some of these methods. Interestingly, while RAM outperforms IMM-JHSE in the car class, IMM-JHSE outperforms RAM in the pedestrian class. While there is a large difference (3.49 HOTA) between IMM-JHSE and the top-ranked method (BiTrack [19]), it should also be noted that IMM-JHSE uses publicly available detections and the corresponding difference in detection accuracy is 3.64%. Furthermore, BiTrack is an offline method. Considering that IMM-JHSE uses only an initial homography estimate and further relies only on 2D detections, it is remarkable that it compares so well to 3D MOT methods (which use regular 3D measurements).
However; in Table 5, which consists of the top three KITTI-pedestrian methods, IMM-JHSE shows the poorest results for the pedestrian class in all metrics except AssA (though it still performs well; it is ranked second on the online KITTI-pedestrian leaderboard — UCMCTrack’s results do not appear there). As will be seen in Table 7 and Table 8 for the MOT17 and MOT20 datasets, IMM-JHSE performs worse with tracks that take up relatively little of the image plane area, particularly if a multitude of such tracks are close to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, suppose multiple tracks are close to one another, and all have ambiguous detections. In that case, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
@@
OC-SORT [11] 76.5 90.3 76.4 54.7 65.1 59.1
UCMCTrack [32] 77.1 90.4 77.2 55.2 67.4 58.0
UCMCTrack+ [32] 74.2 90.2 71.7 54.3 67.2 56.3
@@
Table 5: Results on the KITTI test set with detections obtained from PermaTrack [89].
@@
+PointRCNN [21] 69.99 83.61 69.33 37.81 38.13 44.33
Mono 3D KF [20] 75.47 88.48 77.63 42.87 45.44 46.31
OC-SORT [11] 76.54 90.28 76.39 54.69 65.14 59.08
PermaTrack [89] 78.03 91.33 78.41 48.63 65.98 45.61
@@
RAM [88] 79.53 91.61 80.94 52.71 68.4 52.19
MCTrack (online) [26] 80.78 89.82 84.30
CasTrack [17] 81.00 91.91 84.22
MCTrack (offline) [26] 82.56 91.62 86.64
BiTrack (offline) [19] 82.7 91.55 86.17
@@ 6.3 Results on the MOT datasets
ByteTrack [9] 63.1 77.3 80.3 62.0
C-BIoU [12] 64.1 79.7 81.1 63.7
OC-SORT [11] 63.2 77.5 78.0 63.4
UCMCTrack [32] 64.3 79.0 79.0 64.6
UCMCTrack+ [32] 65.8 81.1 80.5 66.6
@@
MeMOTR [45] 58.8 71.5 72.8 58.4
MOTIP [46] 59.2 71.2 75.5 56.9
Deformable DETR [86]
Table 7: Results on the MOT17 test set. For the tracking-by-detection methods, the detections are obtained from [9].
@@
ByteTrack [9] 61.3 75.2 77.8 59.6
OC-SORT [11] 62.4 76.3 75.7 62.5
UCMCTrack [32] 62.8 77.4 75.5 63.5
UCMCTrack+ [32] 62.8 77.4 75.7 63.4
@@
Table 8: Results on the MOT20 test set with detections obtained from [9].
@@ 7 Conclusion
Furthermore, IMM-JHSE performs remarkably similarly to attention-based methods on the DanceTrack dataset and outperforms them on the MOT17 dataset. Not only does IMM-JHSE bridge the gap between 3D and 2D MOT methods, but also between tracking-by-detection and tracking-by-attention methods.
@@
However, a weak point of IMM-JHSE is that it performs poorly when tracks are very dense, especially in the MOT20 dataset. This is attributed to the leniency afforded by the multitude of motion models in use. Future work may, therefore, improve its performance on data with high track densities.

### v4

@@ 5 Experiments
This section describes the experimental procedure used to find optimal hyperparameters for IMM-JHSE and reports the results of experiments on validation data. The datasets considered during these experiments are the MOT17 (mot17) and DanceTrack (dance) datasets. Results on the test sets are reported in Section 6 and include the MOT20 (mot20) and KITTI (kitti) datasets. All tests and experiments are performed on a single core with a Ryzen 5900X CPU. Inference time is not a major concern in this paper, with optimisation being deferred to future work. Nevertheless, the current state achieves an average of 162.5 frames per second on the MOT17 dataset. Note that since publicly available detections are reused, this is not indicative of the expected inference time in a practical scenario.
For evaluation, the TrackEval library is used (trackeval) to obtain the higher order tracking accuracy (HOTA), association accuracy (AssA), IDF1 and multi-object tracking accuracy (MOTA). It has become standard to report these metrics in the most closely related recent work (ucmc; Yang2022HardSpace), and we refer the reader to the paper which introduces HOTA (hota) for detailed discussions of these metrics, including the use of the Hungarian algorithm. Importantly, (hota) clearly explains the numerous deficiencies of metrics like MOTA and IDF1 that are addressed by HOTA. As such, we consider HOTA the primary evaluation metric in this study. However, since HOTA can be decomposed into separate detection accuracy and association accuracy scores (hota), and publicly available detections are used presently, it is especially appropriate to compare AssA between methods that use the same detections. Since this work utilises publicly available detections, the detection accuracy (DetA) is not reported to avoid distracting from association performance.
5.1 Evaluation metrics
The MOT metrics used in this work are briefly introduced in this subsection.
5.1.1 Multiple object tracking accuracy
MOTA evaluates association accuracy by counting identity switches ($\mathrm{IDSW}$). However, it also takes into account detection false positives (FPs) and false negatives (FNs):
 $\text{ MOTA }=1-\frac{|\mathrm{FN}|+|\mathrm{FP}|+|\mathrm{IDSW}|}{|\mathrm{gtDet}|},$
where $\mathrm{gtDet}$ refers to ground truth detections. Here, $|\cdot|$ denotes the total number of the quantity concerned. In simple terms, MOTA adds the number of false positive and false negative detections to the number of identity switches, normalises the result by the number of ground truth detections and subtracts this from one (hota).
5.1.2 IDF1
IDF1, or identity (ID) F1 score, considers entire track trajectories. Whereas MOTA counts identity switches once each time a track is assigned one ID in one frame and another in the next, IDF1 considers the proportion of overall time that a track is assigned the correct ID over its lifetime. IDF1 is calculated from ID true positives (IDTPs), ID false negatives (IDFNs) and ID false positives (IDFPs) (hota):
$\mathrm{IDF}1=\frac{|\mathrm{IDTP}|}{|\mathrm{IDTP}|+0.5|\mathrm{IDFN}|+0.5|\mathrm{IDFP}|}.$
5.1.3 Higher-order tracking accuracy
HOTA (hota) introduces the concept of true positive associations (TPAs), false negative associations (FNAs) and false positive associations (FPAs), in addition to detection-level true positives (TPs), FPs and FNs. At a localisation threshold $\alpha$, it is calculated as follows:
$\operatorname{HOTA}_{\alpha}=\sqrt{\frac{\sum_{c\in\{\mathrm{TP}\}}\mathcal{A}(c)}{|\mathrm{TP}|+|\mathrm{FN}|+|\mathrm{FP}|}},$
$\mathcal{A}(c)=\frac{|\mathrm{TPA}(c)|}{|\operatorname{TPA}(c)|+|\operatorname{FNA}(c)|+|\mathrm{FPA}(c)|},$
where $\sum_{c\in\{\mathrm{TP}\}}$ indicates the sum over every true positive detection $c$ (hota). The final HOTA score is obtained by averaging $\operatorname{HOTA}_{\alpha}$ over various values of $\alpha$ (hota).
HOTA can also be decomposed into different error types, such as localisation error, detection accuracy, detection precision, detection recall, association accuracy, association precision, and association recall (hota).
5.1.4 Association Accuracy
Association accuracy is calculated as follows:
$\operatorname{AssRe}_{\alpha}=\frac{1}{|\mathrm{TP}|}\sum_{c\in\{\mathrm{TP}\}}\frac{|\mathrm{TPA}(c)|}{|\mathrm{TPA}(c)|+|\operatorname{FNA}(c)|}$
$\operatorname{AssPr}_{\alpha}=\frac{1}{|\mathrm{TP}|}\sum_{c\in\{\mathrm{TP}\}}\frac{|\operatorname{TPA}(c)|}{|\operatorname{TPA}(c)|+|\operatorname{FPA}(c)|}$
$\operatorname{AssA}_{\alpha}=\frac{\operatorname{AssRe}_{\alpha}\cdot\operatorname{AssPr}_{\alpha}}{\operatorname{AssRe}_{\alpha}+\operatorname{AssPr}_{\alpha}-\operatorname{AssRe}_{\alpha}\cdot\operatorname{AssPr}_{\alpha}}.$
Like HOTA, the final AssA score is obtained by averaging $\operatorname{AssA}_{\alpha}$ over various values of $\alpha$. Association precision and recall are also novel contributions of (hota): association recall will be low if a tracker splits a ground truth track into multiple tracks, and association precision will be low when a tracker predicts the same identity for multiple objects. Since IMM-JHSE uses publicly available detections, it is appropriate to consider association accuracy and IDF1 specifically in comparison with other methods that use the same detections.
@@ 5.2 Parameter selection
The largest assumption of IMM-JHSE is that an initial homography estimate is available. For all experiments and test set results, the camera calibration matrices provided in (ucmc) are used – except for the KITTI dataset, which provides its own. Furthermore, the estimates of $\mathbf{\hat{A}}_{t}$ (2) provided by (ucmc) are used for the MOT17 and MOT20 datasets, but for the DanceTrack and KITTI datasets $\mathbf{\hat{A}}_{t}$ is estimated as mentioned in Section 3.
@@ 5.2.1 Filter initialisation
In addition to the parameters above, the covariance matrices with which each filter is initialised must be specified. The initial covariance matrix associated with the ground plane coordinates $\mathbf{x}^{W}_{t}$ is obtained similar to (ucmc):
@@
This paper uses the pattern search algorithm (pattern_search) to find optimum values for the hyperparameters. For each experiment described subsequently, the initial values are set as follows: $\sigma_{x}=5,\sigma_{y}=5,\alpha_{1}=\alpha_{2}=\alpha_{3}=0.5,\Omega=30,b=0,d_{\text{high}}=0.6,d_{\text{low}}=0.5,p_{s,s}=p_{d,d}=p_{W,W}=p_{I,I}=0.9,v=0.5$, and the number of pattern search iterations is restricted to 200. The objective function selected for optimisation is HOTA for the MOT17, MOT20 and KITTI datasets and association accuracy (AssA) for the DanceTrack dataset. Other evaluation metrics include IDF1 and multiple object tracking accuracy (MOTA). The reader is referred to (hota) for details regarding these metrics. In all cases, evaluation is performed with the TrackEval library (trackeval). For the DanceTrack, KITTI and MOT20 datasets, the average of the objective functions of each video in the entire dataset is optimised, while for the MOT17 dataset, the objective function is optimised per video.
The parameters obtained for IMM-JHSE with the pattern search algorithm are given in Table 9. On average, the pattern search algorithm takes 1.2 hours to find the best parameters on each dataset.
@@ 5.3 Ablation study on the DanceTrack dataset
Table 1 reports the results of an ablation study on the DanceTrack validation dataset. The proposed method is focused on handling the situation where targets regularly leave the ground plane. As such, we do not perform ablation studies on other datasets. In light of this, a deficiency of our proposed method is that it is designed for the DanceTrack dataset in particular.
The second column in Table 1 refers to using the dynamic measurement noise estimation method described in Section 4.3. When dynamic measurement noise is not used (as indicated with X), (16) is used to provide an estimate of the measurement noise at each time step. When the dynamic estimation method is not used, HOTA decreases from $63.81$ to $60.80$. Thus, the dynamic noise estimation method contributes significantly to the success of IMM-JHSE.
@@
UCMC (ucmc) X X X X 60.42 62.64
UCMC+ (ucmc) X X X X 59.18 62.52
Table 1: Ablation study on the DanceTrack validation set, with detections obtained from (Cao2022Observation-CentricTracking). The columns denote various components of IMM-JHSE as described in Section 5. UCMC+ refers to the method of (ucmc) with camera motion compensation.
@@ 5.4 Investigating association score mixture methods
Table 2: Validation results on the MOT17 and DanceTrack training sets with detections obtained from (Zhang2022ByteTrack:Box), where the way in which bounding box information is used is varied.
@@ 5.5 Investigation of the chosen objective function
Table 3: Validation results on the MOT17 and DanceTrack training sets with detections obtained from (Zhang2022ByteTrack:Box), where the optimised metric is varied. Note that, for some datasets, the results depend greatly on the metric being optimised.
@@ 6 Results
This section presents the results of IMM-JHSE evaluated on the DanceTrack, MOT17, MOT20 and KITTI test datasets. Note that the source of detections is indicated in the table captions. When comparing to tracking-by-detection methods, the same publicly available detections are used. Thus, enabling a fair comparison between the association methods only. For attention and 3D MOT methods, we refer the reader to the relevant citations for information about the vast variety of detectors and training strategies.
@@ 6.1 Results on DanceTrack
IMM-JHSE uses a single initial homography estimate and maintains the state of the homography projection matrix separately for each track. This strategy may aid association by allowing individual track homography estimates to diverge from the unobserved universal transformation, thus serving as an additional source of disambiguation between tracks. In contrast, attention-based methods are at a disadvantage in this regard since they do not incorporate ground-plane information.
@@
ByteTrack (Zhang2022ByteTrack:Box) 47.3 52.5 89.5 31.4
C-BIoU (Yang2022HardSpace) 60.6 61.6 91.6 45.4
OC-SORT (Cao2022Observation-CentricTracking) 55.1 54.9 92.2 40.4
UCMCTrack (ucmc) 63.4 65.0 88.8 51.1
UCMCTrack+ (ucmc) 63.6 65.0 88.9 51.3
@@
FusionTrack (fusiontrack) 65.3 73.3 90.1 57.5
MeMOTR (Gao2024MeMOTR:Tracking) 66.7 70.6 94 53
MeMOTR (Gao2024MeMOTR:Tracking) 68.5 71.2 89.9 58.4
MOTIP (Gao2024MultiplePrediction) 67.5 72.2 90.3 57.6
MOTIP (Gao2024MultiplePrediction) 70 75.1 91 60.8
Deformable DETR (Zhu2021DeformableDetection)
DAB-Deformable DETR (Liu2022DAB-DETR:DETR)
Table 4: Results on the DanceTrack test set. For the tracking-by-detection methods, the detections are obtained from (Zhang2022ByteTrack:Box).
@@ 6.2 Results on the KITTI dataset
At the time of writing, IMM-JHSE is placed twenty-fourth on the KITTI-car leaderboard [^1] https://www.cvlibs.net/datasets/kitti/eval_tracking_detail.php?result=c3e8939f18cba7b23a2e05a92d4df9e994eaaf91. In addition, all methods that are placed above it are 3D MOT methods that make use of regular 3D measurements, except for RAM (Tokmakov2022ObjectMemory), which uses a self-supervised encoder that optimises a Markov walk along a space-time graph. Table 6 compares IMM-JHSE and some of these methods. Interestingly, while RAM outperforms IMM-JHSE in the car class, IMM-JHSE outperforms RAM in the pedestrian class. While there is a large difference (3.49 HOTA) between IMM-JHSE and the top-ranked method (BiTrack (Huang2024BiTrack:Data)), it should also be noted that IMM-JHSE uses publicly available detections, and the corresponding difference in detection accuracy is 3.64%. Furthermore, BiTrack is an offline method. Considering that IMM-JHSE uses only an initial homography estimate and further relies only on 2D detections, it is remarkable that it compares so well to 3D MOT methods (which use regular 3D measurements). This implies that the additional disambiguating information afforded by 3D methods is closely matched by the expanded track state maintained in IMM-JHSE and its novel association method. Practically, the use of IMM-JHSE foregoes processing 3D measurements in every frame, although this comes at the cost of a more complex state and association model.
In Table 5, which consists of the top three KITTI-pedestrian methods, IMM-JHSE shows the poorest results for the pedestrian class in all metrics except AssA (though it still performs well; it is ranked second on the online KITTI-pedestrian leaderboard — UCMCTrack’s results do not appear there). As will be seen in Table 7 and Table 8 for the MOT17 and MOT20 datasets, IMM-JHSE performs worse with tracks that take up relatively little of the image plane area, particularly if a multitude of such tracks are close to one another. In other words, the merits of IMM-JHSE are more evident when tracking relatively large (in terms of bounding box area) objects that move irregularly or with high acceleration or velocity in the image plane rather than in highly cluttered situations. This may be due to the second association stage, which is designed to account for ambiguous detections that may diverge from that expected by the predicted ground plane state. Yet, suppose multiple tracks are close to one another, and all have ambiguous detections. In that case, it is conceivable that they may be confused with one another since they are likely to have high BIoU scores with detections that belong to other tracks.
@@
OC-SORT (Cao2022Observation-CentricTracking) 76.5 90.3 76.4 54.7 65.1 59.1
UCMCTrack (ucmc) 77.1 90.4 77.2 55.2 67.4 58.0
UCMCTrack+ (ucmc) 74.2 90.2 71.7 54.3 67.2 56.3
@@
Table 5: Results on the KITTI test set with detections obtained from PermaTrack (permatrack).
@@
+PointRCNN (Weng20203DMetrics) 69.99 83.61 69.33 37.81 38.13 44.33
Mono 3D KF (Reich2021MonocularTracks) 75.47 88.48 77.63 42.87 45.44 46.31
OC-SORT (Cao2022Observation-CentricTracking) 76.54 90.28 76.39 54.69 65.14 59.08
PermaTrack (permatrack) 78.03 91.33 78.41 48.63 65.98 45.61
@@
RAM (Tokmakov2022ObjectMemory) 79.53 91.61 80.94 52.71 68.4 52.19
MCTrack (online) (Wang2024MCTrack:Driving) 80.78 89.82 84.30
CasTrack (Wu2022CasA:Clouds) 81.00 91.91 84.22
MCTrack (offline) (Wang2024MCTrack:Driving) 82.56 91.62 86.64
BiTrack (offline) (Huang2024BiTrack:Data) 82.7 91.55 86.17
@@ 6.3 Results on the MOT datasets
ByteTrack (Zhang2022ByteTrack:Box) 63.1 77.3 80.3 62.0
C-BIoU (Yang2022HardSpace) 64.1 79.7 81.1 63.7
OC-SORT (Cao2022Observation-CentricTracking) 63.2 77.5 78.0 63.4
UCMCTrack (ucmc) 64.3 79.0 79.0 64.6
UCMCTrack+ (ucmc) 65.8 81.1 80.5 66.6
@@
MeMOTR (Gao2024MeMOTR:Tracking) 58.8 71.5 72.8 58.4
MOTIP (Gao2024MultiplePrediction) 59.2 71.2 75.5 56.9
Deformable DETR (Zhu2021DeformableDetection)
Table 7: Results on the MOT17 test set. For the tracking-by-detection methods, the detections are obtained from (Zhang2022ByteTrack:Box).
@@
ByteTrack (Zhang2022ByteTrack:Box) 61.3 75.2 77.8 59.6
OC-SORT (Cao2022Observation-CentricTracking) 62.4 76.3 75.7 62.5
UCMCTrack (ucmc) 62.8 77.4 75.5 63.5
UCMCTrack+ (ucmc) 62.8 77.4 75.7 63.4
@@
Table 8: Results on the MOT20 test set with detections obtained from (Zhang2022ByteTrack:Box).
@@ 7 Conclusion
Furthermore, IMM-JHSE performs remarkably similarly to attention-based methods on the DanceTrack dataset and outperforms them on the MOT17 dataset. Not only does IMM-JHSE bridge the gap between 3D and 2D MOT methods, but also between tracking-by-detection and tracking-by-attention methods. This could indicate that IMM-JHSE isolates some of the advantages afforded by 3D and attention-based MOT methods. Particularly, IMM-JHSE seems to indicate that the disambiguating information between tracks afforded by maintaining separate joint homography estimates (based only on an initial measurement) rivals that of 3D and attention-based methods. Crucially, IMM-JHSE does not guarantee an accurate global homography estimate, but focuses on maintaining an accurate projection for each track separately. Because IMM-JHSE requires only an initial homography estimate, the complexity of handling 3D measurements in every frame (or a context window for attention-based methods) can be foregone. However, this comes at the cost of maintaining a more complex state and association model.
@@
However, a weak point of IMM-JHSE is that it performs poorly when tracks are very dense, especially in the MOT20 dataset. This is attributed to the leniency afforded by the multitude of motion models in use and the fact that IMM-JHSE is designed primarily with the DanceTrack dataset in mind. Future work may, therefore, improve its performance on data with high track densities and generalisation to other datasets. It is also possible that results may be significantly improved with the use of a custom detector module.